OpenStudy (anonymous):

Write an expression in simplest form for the volume of a cube with a side of 11x6 units. a. 121x12 b. 1331x18 c.121x36 d.1331x12

OpenStudy (anonymous):

@Reaper534

OpenStudy (anonymous):

\[V=a^3\]

OpenStudy (anonymous):

Thats the volume of a cube ok

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Where _a_ is the length of the side. If the side is 1^6 = 1 then the volume is 1^3 which is 1 too. If the side is 11*x^6 then the volume is (11*x^6)^3 = 1331*x^18 What ever the length of the side is, raise it to the power of 3 :)

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

So your answer is B

OpenStudy (anonymous):

Did you get what i wrote up there lol?